I'm having trouble with solving systems. I did this before, at the beginning of the year, but I completely forget how to do it. Here's the system:
y=7x-9
y=4x+6

Answer by smartdude17(30)

Well for a system you start like this:
0=7x-9 G,lety=0
0=4x+6
4x+6=7x-9 Merge by Sub
15/3=3x/3 S
5=x Solution
4(5)+6=7(5)-9 Sub
26=26